# Many-Body Localization in the Age of Classical Computing

### v2

@@ IX.2.4 Large-deviation approaches to resonances
Understanding many-body resonances as eigenstates that are superpositions of a small number of basis states reinforces the significance of the Fock space perspective on MBL [379, 532, 641]. To enhance the effects of rare configurations corresponding to long-range resonances, Ref. [79] introduces a biased sampling scheme that magnifies the impact of the tails of the probability distribution of the transmission amplitudes.
Considering the propagator $G_{j,0}=\braket{j}{H^{-1}}{0}$ at energy $E=0$ between a fixed state |$\ket{0}$, and states |$\ket{j}$ of the eigenbasis of $S^{z}_{j}$ operators,  [79] examine the quantity
@@
where ${\cal J}_{n}$ denotes an exponentially large set of basis states differing by a fixed number of spin flips $n$ from the state |$\ket{0}$. The function ${\cal T}_{0}(\beta)$ is analogous to a statistical mechanics model partition function. For $\beta=2$, the quantity ${\cal T}_{0}(2)$, in terms of the single-particle hopping on the Fock space graph, can be interpreted as being proportional to the total Landauer transmission corresponding to the situation when the particles are injected at the point |$\ket{0}$ and extracted at the node |$\ket{j}$ for $j\in\cal{J}$. The basic idea is, however, to vary the auxiliary parameter $\beta$, tuning the relative contribution from members |$\ket{j}$ of $\cal{J}$ that form strong resonances with the spin configuration |$\ket{0}$, corresponding to significant values of $G_{j,0}$.
@@
Analyzing dependence of a free energy $\phi(\beta)=\frac{1}{\beta n}\ln\braket{{\cal T}_{0}(\beta)}$ of the value of $\beta$[79] distinguish four regimes of disorder strength characterized by different behavior of the propagator $G_{j,0}$. At disorder strength $W<W_{\mathrm{ergo}}$, all self-avoiding paths on the Fock space graph between |$\ket{0}$ and the states |$\ket{j}$ contribute to the dynamics. The system resembles a fully connected quantum dot in that regime, and the level statistics approach the RMT predictions. At larger disorder strengths $W_{\mathrm{ergo}}<W<W_{g}$, the number of self-avoiding paths contributing to the dynamics grows exponentially with $n$ but only a sub-extensive fraction of paths contributes, and the level statistics is no longer following the RMT prediction. For $W_{g}<W<W_{c}$, the state |$\ket{0}$ hybridizes only with a few, $O(1)$ states |$\ket{j}$ far away in the Hilbert space. Those rare many-body resonances are unraveled thanks to tuning the value of $\beta$. Finally, at $W>W_{c}$ the probability to evolve away from the initial state |$\ket{0}$ goes to zero exponentially with $n$ even when the system-wide resonances are accounted for. Interestingly, Ref. [79] reports that the drift of the disorder strength $W_{c}$ with system size $L$ is very mild. Moreover, the same approach based on (86) to Anderson model on Cayley tree reproduces the known critical disorder strength $W_{c}$ [472, 657] with a good accuracy.
@@
Interesting directions for future studies involve linking the four disorder regimes found in [79] with properties of rare thermal inclusions and understanding the relation of the four disorder regimes to the many-body resonance probes considered in [427]. Finally, it would be interesting to understand whether the biased sampling method of [79] can be formulated for quantities that can be probed by the time evolution of many-body systems, increasing the relevance of the method both from theoretical and experimental standpoints. The described future directions may shed new light on the status of MBL in disordered spin chains.
@@ X Possible extensions
In this section, we briefly describe investigations of MBL in many-body systems different than the disordered spin chains on which we focused so far. We start with disordered one-dimensional Hamiltonians with short-range interactions, which share the gross features characterizing ergodicity breaking in the XXZ spin chain, but also host a number of distinguishing features. We then proceed by concisely outlining the studies of MBL in systems with quasiperiodic potential, in disordered Floquet systems, in systems with the presence of long-range interaction, in systems beyond one spatial dimension, and in systems without disorder.
@@ X.1 Other models of MBL
where $J_{i}$, $h_{i}$ and $\gamma_{i}$ are random couplings selected from certain distributions. The clean system with non-vanishing couplings $J_{i}=J$, $h_{i}=h$, and $\gamma_{i}=\gamma$ is an example of non-intergrable ergodic system with ballistic spreading of entanglement and diffusive energy transport [324]. One choice of the couplings is $\gamma_{i}=1$, $J_{i}\in[0.8,1.2]$, and $h_{i}\in[-W,W]$. In that case, tuning the disorder strength $W$ leads to a crossover between the ETH and MBL regimes [1, 164]. The finite size drifts at the crossover are analogous to the case of the disordered XXZ and $J_{1}$-$J_{2}$ models described in Sec. VI.1. For instance, as shown in [596], the crossing point can be approximated as $W^{*}(L)\propto 1/L$, while the boundary of the ergodic regime drifts approximately linearly with the system size $L$, $W^{T}(L)\propto L$.
The disordered transverse field Ising model, in contrast to the disordered XXZ spin chain (7), has no $U(1)$ symmetry. This leads to a larger dimension of the Hilbert space, $\mathcal{N}=2^{L}$, at a fixed system size $L$. However, the geometry of the Fock-space graph is significantly simpler than in the XXZ spin chain. Each of the states of computational basis (i.e., the eigenbasis of the $\hat{S}^{z}_{i}$ operators) is linked by the Hamiltonian to exactly $L$ states differing by a single spin flip, simplifying the Fock-space perspective on MBL in TFIM [536, 142].
Hamiltonian obtained as $\hat{H}^{\prime}=\hat{H}_{\mathrm{TFIM}}+\sum_{i}g_{i}\hat{S}_{i}^{z}\hat{S}_{i+2}^{z}$ has a $Z_{2}$ symmetry given by the operator $\prod_{j=1}^{L}\hat{S}_{i}^{x}$ provided that $h_{i}=0$ is set. Due to the breakdown of the ergodicity in the MBL regime, the $Z_{2}$ symmetry may be broken not only in the ground state, but also in highly excited eigenstates, as proposed by [289] [^36] Analogously, the Anderson localization in non-interacting models leads to breakdown of the $Z_{2}$ symmetry at all energy densities as review in [350].. Numerical evidence for MBL regimes with paramagnetic and spin glass order at all energy densities was found in [326]. Notably, the ergodicity was suggested [433, 544] to be restored at crossovers between the different MBL regimes, and the wide extension of the intervening ergodic regime was verified by [352] [^37] Tensor network investigations of the ground state of the disordered quanum Ising chain find that the single-particle localization is stable against finite interactions [123]. This suggest presence of a crossover to ergodic states at finite energy density above the ground state.. Signatures of topological MBL phases [116] were observed in a variant of the disordered Ising spin chain in [352].
@@
$\hat{H}_{\mathrm{FH}}=\sum_{i,\sigma=\uparrow,\downarrow}J(\hat{c}^{{\dagger}}_{i\sigma}\hat{c}_{i+1,\sigma}+h.c)+\sum_{i}Un_{i\uparrow}n_{i\downarrow}+\sum_{i}\mu_{i}\hat{n}_{i},$ (88)
@@
where $c^{{\dagger}}_{i,\sigma},\hat{c}_{i\sigma}$ are creation and annihilation operators for fermions of spin $\sigma=\uparrow,\downarrow$, $J=-1$ and $U$ are tunneling and interaction amplitudes, respectively, $\mu_{i}\in[-W;W]$ is a random on-site potential, and the sum over $i$ extends from $1$ to the system size $L$ Defining $\hat{n}_{i,\sigma}=\hat{c}^{{\dagger}}_{i,\sigma}\hat{c}_{i\sigma}$, we observe that $\hat{H}_{\mathrm{FH}}$ separately conserves the total number of fermions of spin up, $\hat{N}_{\uparrow}=\sum_{i}\hat{n}_{i,\uparrow}$, and spin down, $\hat{N}_{\downarrow}=\sum_{i}\hat{n}_{i,\downarrow}$. An optical lattice realizing the Fermi-Hubbard model with a quasiperiodic potential playing the role of the disorder $\mu_{i}$, see Sec. X.2, allowed the first observation of signatures of the MBL regime [561].
@@
The on-site Hilbert space of the Fermi-Hubbard is four dimensional, which leads to richer physics than in the disordered spin-1/2 chains. The Hamiltonian (88) possesses additional parity symmetry, and SU(2) symmetry which may be broken by an addition of weak magnetic field and chemical potential terms at the edges of the chain, enabling observation of the ETH-MBL crossover with the average gap ratio [422]. Moreover, the disorder $\mu_{i}$ in Eq. (88) is a charge disorder, i.e., the fermions of different spins experience the same value of $\mu_{i}$. This leads to distinct dynamical behaviors of the charge degrees of freedom, $\hat{n}_{i}=\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow}$, and spin degrees of freedom, $\hat{s}^{z}_{i}=(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})/2$. It was proposed [500] that spin degrees of freedom remain ergodic, while charge degrees of freedom are in the MBL regime. This scenario has been further investigated numerically in [726, 340, 511, 360, 620]. Another possibility is to include a spin disorder upon replacing the term $\sum_{i}\mu_{i}\hat{n}_{i}$ in $\hat{H}_{\mathrm{FH}}$ by $\sum_{i}\mu_{i}\hat{s}^{z}_{i}$. In that case, the model has an $\eta$-pairing symmetry [713] and it is possible to construct a large number of eigenstates of $\hat{H}_{\mathrm{FH}}$, which are characterized both by area-law entanglement entropy, as well as eigenstates with entanglement scaling logarithmically with system size $L$. Remarkably, the $\eta$-pairing symmetry allows to numerically construct highly excited eigenstates of systems comprising up to few hundred lattice sites [722].
@@
From the point of view of the currently understood trends towards ergodicity, observed in the disordered spin chains, it would be interesting to reconsider the case of charge disorder and understand to which degree the ergodic spin degrees of freedom may serve as a bath for the localized charge degrees of freedom. One step in this direction was done by [345] and followed by [92, 93] which consider a single mobile impurity, $\hat{N}_{\uparrow}=1$ in contact with $\hat{N}_{\downarrow}=L/3$ particles of different species in Anderson localized state [^38] Refs. [92, 93] investigate a system of $L/3$ hard-core bosons interacting with a single boson from different species. The latter model can be mapped to Fermi-Hubbard Hamiltonian upon a generalization of Jordan-Wigner transformation considered in [584, 509].. In that case, the on-site potential term in $\sum_{i}\mu_{i}\hat{n}_{i}$ is replaced by $\sum_{i}\mu_{i}\hat{n}_{i,\downarrow}$, and the mobile impurity playes a role of an $L$-level heat bath. The Anderson localization regime is argued to remain stable in the regime of strong interaction between the two species of particles. In that case, the singe impurity becomes immobile and the interactions effectively turn the Anderson insulator into an MBL system with non-trivial entanglement entropy dynamics. Subsequent [590] numerical simulations indicate that there is a wide regime of parameters in which the impurity spreads subdiffusively and induces a gradual delocalization of the Anderson insulator. Nevertheless, at sufficiently large interaction and disorder strength the dynamics remains non-ergodic at all accessible system sizes and time scales [91].
@@
The Bose-Hubbard model [225] is a fundamental model describing bosons hopping on a lattice, and hosting a superfluid to Mott insulator phase transition [207] realized in optical lattices [249]. The Bose-Hubbard Hamiltonian reads
@@
where $\hat{a}^{{\dagger}}_{i},\hat{a}_{i}$ are bosonic creation and annihilation operators, the tunneling amplitude $J=-1$ sets the energy scale, $U>0$ is the interaction strength, the chemical potential $\mu_{i}$ is distributed uniformly in an interval $[-W;W]$, and $\langle i,j\rangle$ is a sum over neighboring sites of 1D lattice comprising $L$ sites. Defining the bosonic number operator as $\hat{n}_{i}=\hat{a}^{{\dagger}}_{i}\hat{a}_{i}$, we observe that $\hat{H}_{BH}$ conserves the total number of bosons $\hat{N}=\sum_{i}\hat{n}_{i}$.
Increase of the disorder strength $W$ in the disordered Bose-Hubbard model (89) leads to a crossover between the ETH and MBL regimes [601, 454, 276]. The main difference between the Bose-Hubbard model and the disordered spin-1/2 chain is that the number of the bosons $\hat{n}_{i}$ on a site $i$ is not restricted in the former model. At a fixed density of particles $\nu=N/L$, the states with atypically large particle occupation numbers lie at large energy densities. The density of states at these large energy densities is small, which leads to an inverted many-body mobility edge observed numerically in the disordered Bose-Hubbard model [601]. The question whether this inverted many-body mobility edge persists in the asymptotic limit $L,t\to\infty$ remains open.
@@
Considering the limit of strong interactions $J_{i}\to\infty$ in the TFIM (87) and reducing the description to the lowest energy subspace, we arrive at the (generalized) PXP Hamiltonian
$\hat{H}_{\mathrm{PXP}}=\sum_{i=1}^{L}P^{\alpha}_{i}S^{x}_{i}P^{\alpha}_{i+1+\alpha}+\sum_{i=1}^{L}h_{i}S^{z}_{i},$ (90)
@@
where the projectors $P^{\alpha}_{i}=\prod_{j=i-\alpha}^{i-1}(1/2-S^{z}_{j})$ assure that the dynamics is confined to a constrained subspace of the full Hilbert space, in which sites with state $\ket{\uparrow}$ are separated by $\alpha$ sites. The limit of strong interactions in the TFIM corresponds to the blockade radius $\alpha=1$, however, this model can be also considered for a more general $\alpha>1$. In Eq. (90), the values of $h_{i}$ are independent and uniformly distributed random variables in the interval $[-W/2,W/2]$ with $W$ being the disorder strength, and periodic boundary conditions (PBC) $\vec{S}_{L+i}\equiv\vec{S}_{i}$ are assumed.
@@
For $\alpha=1$, the model in Eq. (90) is known as the PXP model. In the clean limit, $h_{i}=0$, the PXP Hamiltonian is a paradigmatic model of quantum many-body scars [364, 669, 575] describing the experiments with arrays of Rydberg atoms [61, 94], and can be mapped onto a U(1) gauge theory [629]. In presence of disorder, $W>0$, the PXP model hosts a crossover between the ETH and MBL regimes [120]. However, the system size drifts at this crossover are significantly more pronounced than in the disordered XXZ spin chain, and both disorder strengths $W^{T}(L)$ as well as $W^{*}(L)$ shift linearly with $L$ [594].
@@
The Hilbert space dimension of the PXP model (90) scales as $\mathcal{N}\propto(\Phi_{\alpha})^{L}$, where $\Phi_{\alpha}<2$ decreases monotonously with the growth of the blockade radius $\alpha$. Hence, for larger $\alpha$ access to larger system sizes $L$ with exact diagonalization methods is possible. Notably, at $\alpha>1$, the linear trends $W^{T}(L)\propto L$ and $W^{*}(L)\propto L$ are even more pronounced, suggesting that the PXP models (90) remain ergodic in the asymptotic limit. The nature of the kinetic constraints in $\hat{H}_{\mathrm{PXP}}$ is one factor responsible for this behavior. Indeed, the effective radius of the Fock space graph is much smaller in the PXP model than in the TFIM, enhancing the trends towards ergodicity in the former model [594]. Notably, Ref. [537] reported an opposite trend of enhancing the MBL regime due to the interplay between interaction and kinetic constraints that are of different type than those considered here.
@@
The milder growth of the Hilbert space with increase of the number of lattice sites in constrained spin models has been also used in investigations of 2D models [648, 482]. Moreover, the local constraint resemble, to some degree, the local Gauss law in lattice gauge theories. The latter have been shown to host a pronounced region in which the ergodicity is broken. The boundary of this region, i.e., the disorder strength $W^{T}(L)$, is nearly independent of system size $L$ in the 1D Schwinger model with disordered background charges [233].

### v3

@@ IX.2.4 Large-deviation approaches to resonances
Understanding many-body resonances as eigenstates that are superpositions of a small number of basis states reinforces the significance of the Fock space perspective on MBL [415, 578, 692]. To enhance the effects of rare configurations corresponding to long-range resonances, Ref. [82] introduced a biased sampling scheme that magnifies the impact of the tails of the probability distribution of the transmission amplitudes.
Considering the propagator $G_{j,0}=\braket{j}{H^{-1}}{0}$ at energy $E=0$ between a fixed state |$\ket{0}$, and states |$\ket{j}$ of the eigenbasis of $\hat{S}^{z}_{j}$ operators, [82] examined the quantity
@@
where ${\cal J}_{n}$ denotes an exponentially large set of basis states differing by a fixed number of spin flips $n$ from the state |$\ket{0}$. The function ${\cal T}_{0}(\beta)$ is analogous to a partition function of statistical mechanics model defined on a Cayley tree [176], with correlated energies $\omega_{j,0}=-\ln(G_{j,0})$ and the inverse temperature $\beta$. This analogy motivates the quantitative investigations of [82]. For $\beta=2$, the quantity ${\cal T}_{0}(2)$, in terms of the single-particle hopping on the Fock space graph, can be interpreted as being proportional to the total Landauer transmission corresponding to the situation when the particles are injected at the point |$\ket{0}$ and extracted at the node |$\ket{j}$ for $j\in\cal{J}$. The basic idea is, however, to vary the auxiliary parameter $\beta$, tuning the relative contribution from members |$\ket{j}$ of $\cal{J}$ that form strong resonances with the spin configuration |$\ket{0}$, corresponding to significant values of $G_{j,0}$.
@@
Analyzing dependence of a free energy $\phi(\beta)=\frac{1}{\beta n}\ln\braket{{\cal T}_{0}(\beta)}$ of the value of $\beta$[82] distinguished four regimes of disorder strength characterized by different behavior of the propagator $G_{j,0}$. At disorder strength $W<W_{\mathrm{ergo}}$, all self-avoiding paths on the Fock space graph between |$\ket{0}$ and the states |$\ket{j}$ contribute to the dynamics. The system resembles a fully connected quantum dot in that regime, and the level statistics approach the RMT predictions. At larger disorder strengths $W_{\mathrm{ergo}}<W<W_{g}$, the number of self-avoiding paths contributing to the dynamics grows exponentially with $n$ but only a sub-extensive fraction of paths contributes, and the level statistics is no longer follows the RMT prediction. For $W_{g}<W<W_{c}$, the state |$\ket{0}$ hybridizes only with a few, $O(1)$ states |$\ket{j}$ far away in the Hilbert space. Those rare many-body resonances are unraveled thanks to tuning the value of $\beta$. Finally, at $W>W_{c}$ the probability to evolve away from the initial state |$\ket{0}$ goes to zero exponentially with $n$ even when the system-wide resonances are accounted for. Interestingly, Ref. [82] reported that the drift of the disorder strength $W_{c}$ with system size $L$ is very mild. Moreover, the same approach based on (86) to the Anderson model on a Cayley tree reproduces the known critical disorder strength $W_{c}$ [510, 709] with a good accuracy.
@@
Interesting directions for future studies involve linking the four disorder regimes found in [82] with properties of rare thermal inclusions, and understanding the relation of the four disorder regimes to the many-body resonance probes considered in [463]. Finally, it would be interesting to understand whether the biased sampling method of [82] can be formulated for quantities that can be probed by the time evolution of many-body systems, increasing the relevance of the method both from theoretical and experimental standpoints. The described future directions may shed new light on the status of MBL in the disordered spin chains.
@@ X Possible extensions
In this Section, we briefly describe investigations of MBL in many-body systems different than the disordered spin chains on which we focused so far. We start with disordered one-dimensional Hamiltonians with short-range interactions, which share the gross features characterizing ergodicity breaking in the disordered XXZ spin chain, but also host a number of distinguishing features. We then proceed by concisely outlining the studies of MBL in systems with quasiperiodic potential, in disordered Floquet systems, in systems with the presence of long-range interaction, in systems beyond one spatial dimension, and in systems without disorder.
@@ X.1 Other models of MBL
where $J_{i}$, $h_{i}$ and $\gamma_{i}$ are random couplings selected from certain distributions. The clean system with non-vanishing couplings $J_{i}=J$, $h_{i}=h$, and $\gamma_{i}=\gamma$ is an example of non-intergrable ergodic system with ballistic spreading of entanglement and diffusive energy transport [350]. One choice of the couplings is $\gamma_{i}=1$, $J_{i}\in[0.8,1.2]$, and $h_{i}\in[-W,W]$. In that case, tuning the disorder strength $W$ leads to a crossover between the ETH and MBL regimes [1, 171]. The finite size drifts at the crossover appear analogous to the case of the disordered XXZ and $J_{1}$-$J_{2}$ models described in Sec. VI.1. For instance, as shown in [645], the crossing point shifts to larger disorder strengths $W$ sub-linearly with the increase of $L$ and can be approximated as $W^{*}(L)\propto 1/L$. The boundary of the ergodic regime in the disordered TFIM drifts approximately linearly with the system size $L$, $W^{T}(L)\propto L$. The Hamiltonian (87) belongs to the class of models considered in [573].
The disordered transverse field Ising model, in contrast to the disordered XXZ spin chain (7), has no $U(1)$ symmetry. This leads to a larger dimension of the Hilbert space, $\mathcal{N}=2^{L}$, at a fixed system size $L$. However, the geometry of the Fock-space graph is significantly simpler than in the disordered XXZ spin chain. Each of the states in the computational basis (i.e., the eigenbasis of the $\hat{S}^{z}_{i}$ operators) is linked by the Hamiltonian to exactly $L$ states differing by a single spin flip, simplifying the Fock-space perspective on MBL in the TFIM [582, 147].
Hamiltonian obtained as $\hat{H}^{\prime}=\hat{H}_{\mathrm{TFIM}}+\sum_{i}g_{i}\hat{S}_{i}^{z}\hat{S}_{i+2}^{z}$ has a $Z_{2}$ symmetry given by the operator $\prod_{j=1}^{L}\hat{S}_{i}^{x}$, provided that $h_{i}=0$ is set. Due to the breakdown of the ergodicity in the MBL regime, the $Z_{2}$ symmetry may be broken not only in the ground state, but also in highly excited eigenstates, as proposed by [311] [^46] Analogously, the Anderson localization in non-interacting models leads to breakdown of the $Z_{2}$ symmetry at all energy densities as reviewed in [379].. Numerical evidence for the MBL regimes with paramagnetic and spin glass order at all energy densities was found in [353]. Notably, the ergodicity was suggested [470, 590] to be restored at crossovers between the different MBL regimes, and the wide extension of the intervening ergodic regime was verified by [381] [^47] Tensor network investigations of the ground state of the disordered quanum Ising chain find that the single-particle localization is stable against finite interactions [129]. This suggest presence of a crossover to ergodic states at finite energy density above the ground state.. Signatures of topological MBL phases [121] were observed in a variant of the disordered Ising spin chain in [381].
@@
$\hat{H}_{\mathrm{FH}}=\sum_{i,\sigma=\uparrow,\downarrow}J(\hat{c}^{{\dagger}}_{i\sigma}\hat{c}_{i+1,\sigma}+h.c)+\sum_{i}U\hat{n}_{i\uparrow}\hat{n}_{i\downarrow}+\sum_{i}\mu_{i}\hat{n}_{i},$ (88)
@@
where $c^{{\dagger}}_{i,\sigma},\hat{c}_{i\sigma}$ are creation and annihilation operators for fermions of spin $\sigma=\uparrow,\downarrow$, $J=-1$ and $U$ are tunneling and interaction amplitudes, respectively, $\mu_{i}\in[-W;W]$ is a random on-site potential, and the sum over $i$ extends from $1$ to the system size $L$ Defining $\hat{n}_{i,\sigma}=\hat{c}^{{\dagger}}_{i,\sigma}\hat{c}_{i\sigma}$, we observe that $\hat{H}_{\mathrm{FH}}$ separately conserves the total number of fermions of spin up, $\hat{N}_{\uparrow}=\sum_{i}\hat{n}_{i,\uparrow}$, and spin down, $\hat{N}_{\downarrow}=\sum_{i}\hat{n}_{i,\downarrow}$. An optical lattice realizing the Fermi-Hubbard model with a quasiperiodic potential playing the role of the disorder $\mu_{i}$, see Sec. X.2, allowed for the first observation of signatures of the MBL regime [608].
@@
The on-site Hilbert space of the Fermi-Hubbard is four dimensional, which leads to richer physics than in the disordered spin-1/2 chains. The Hamiltonian (88) possesses an additional parity symmetry, and the SU(2) symmetry that may be broken by an addition of weak magnetic field and chemical potential terms at the edges of the chain, enabling observation of the ETH-MBL crossover with the average gap ratio [458]. Moreover, the disorder $\mu_{i}$ in Eq. (88) is a charge disorder, i.e., the fermions of different spins experience the same value of $\mu_{i}$. This leads to distinct dynamical behaviors of the charge degrees of freedom, $\hat{n}_{i}=\hat{n}_{i,\uparrow}+\hat{n}_{i,\downarrow}$, and spin degrees of freedom, $\hat{s}^{z}_{i}=(\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow})/2$. It was proposed [541] that spin degrees of freedom remain ergodic, while charge degrees of freedom are in the MBL regime. This scenario has been further investigated numerically in [782, 369, 552, 390, 670, 564]. Another possibility is to include a spin disorder upon replacing the term $\sum_{i}\mu_{i}\hat{n}_{i}$ in $\hat{H}_{\mathrm{FH}}$ by $\sum_{i}\mu_{i}\hat{s}^{z}_{i}$. In that case, the model has an $\eta$-pairing symmetry [769] and it is possible to construct a large number of eigenstates of $\hat{H}_{\mathrm{FH}}$, which are characterized both by area-law entanglement entropy, as well as eigenstates with the entanglement entropy scaling logarithmically with system size $L$. Remarkably, the $\eta$-pairing symmetry allows to numerically construct highly excited eigenstates of systems comprising up to few hundred lattice sites [778].
@@
From the point of view of the currently understood trends towards ergodicity, observed in the disordered spin chains, it would be interesting to reconsider the case of charge disorder and understand to which degree the ergodic spin degrees of freedom may serve as a bath for the localized charge degrees of freedom. One step in this direction was done by [374] and followed by [95, 96] which consider a single mobile impurity, $\hat{N}_{\uparrow}=1$, in contact with $\hat{N}_{\downarrow}=L/3$ particles of different species in Anderson localized state [^48] Refs. [95, 96] investigated a system of $L/3$ hard-core bosons interacting with a single boson from different species. The latter model can be mapped to Fermi-Hubbard Hamiltonian upon a generalization of Jordan-Wigner transformation considered in [631, 550].. In that case, the on-site potential term in $\sum_{i}\mu_{i}\hat{n}_{i}$ is replaced by $\sum_{i}\mu_{i}\hat{n}_{i,\downarrow}$, and the mobile impurity playes a role of an $L$-level heat bath. The Anderson localization regime is argued to remain stable in the regime of strong interaction between the two species of particles. In that case, the single impurity becomes immobile and the interactions effectively turn the Anderson insulator into an MBL system with non-trivial entanglement entropy dynamics. Subsequent [639] numerical simulations indicate that there is a wide regime of parameters in which the impurity spreads subdiffusively and induces a gradual delocalization of the Anderson insulator. Nevertheless, at sufficiently large interaction and disorder strength the dynamics remain non-ergodic at all accessible system sizes and time scales [94].
@@
The Bose-Hubbard model [241] is a fundamental model describing bosons hopping on a lattice, and hosting a superfluid to Mott insulator ground state phase transition [220], which was realized with ultracold bosons in optical lattices [265]. The Bose-Hubbard Hamiltonian reads
@@
where $\hat{a}^{{\dagger}}_{i},\hat{a}_{i}$ are bosonic creation and annihilation operators, the tunneling amplitude $J=-1$ sets the energy scale, $U>0$ is the interaction strength, the chemical potential $\mu_{i}$ is distributed uniformly in an interval $[-W;W]$, and $\langle i,j\rangle$ is a sum over the neighboring sites of 1D lattice comprising $L$ sites. Defining the bosonic number operator as $\hat{n}_{i}=\hat{a}^{{\dagger}}_{i}\hat{a}_{i}$, we observe that $\hat{H}_{BH}$ conserves the total number of bosons $\hat{N}=\sum_{i}\hat{n}_{i}$.
Increase of the disorder strength $W$ in the disordered Bose-Hubbard model (89) leads to a crossover between the ETH and MBL regimes [650, 492, 295]. The main difference between the Bose-Hubbard model and the disordered spin-1/2 chain is that the number of the bosons $n_{i}$ on a site $i$ is not restricted in the former model. At a fixed density of particles $\nu=N/L$, the states with atypically large particle occupation numbers lie at large energy densities. The density of states at these large energy densities is small, which leads to an inverted many-body mobility edge observed numerically in the disordered Bose-Hubbard model [650], see also [365]. The question whether this inverted many-body mobility edge persists in the asymptotic limit $L,t\to\infty$ remains open.
@@
Replacing the term $\sum_{i=1}^{L}J_{i}\hat{S}_{i}^{z}\hat{S}_{i+1}^{z}$ by $\sum_{i=1}^{L}J_{i}(\hat{S}_{i}^{z}+1/2)(\hat{S}_{i+1}^{z}+1/2)$ in the TFIM (87), taking the limit of strong interactions, ${J_{i}\to\infty}$, and reducing the description to the lowest energy subspace, we arrive at the (generalized) PXP Hamiltonian,
$\hat{H}_{\mathrm{PXP}}=\sum_{i=1}^{L}\hat{P}^{\alpha}_{i}\hat{S}^{x}_{i}\hat{P}^{\alpha}_{i+1+\alpha}+\sum_{i=1}^{L}h_{i}\hat{S}^{z}_{i},$ (90)
@@
where the projectors $\hat{P}^{\alpha}_{i}=\prod_{j=i-\alpha}^{i-1}(1/2-\hat{S}^{z}_{j})$ assure that the dynamics are confined to a constrained subspace of the full Hilbert space, in which sites with state $\ket{\uparrow}$ are separated by $\alpha$ sites. The limit of strong interactions in the TFIM corresponds to the blockade radius $\alpha=1$, however, this model can be also considered for a more general $\alpha>1$. In Eq. (90), the values of $h_{i}$ are independent and uniformly distributed random variables in the interval $[-W/2,W/2]$ with $W$ being the disorder strength, and periodic boundary conditions (PBC) $\vec{S}_{L+i}\equiv\vec{S}_{i}$ are assumed.
@@
For $\alpha=1$, the model in Eq. (90) is known as the PXP model. In the clean limit, $h_{i}=0$, the PXP Hamiltonian is a paradigmatic model of quantum many-body scars [396, 723, 622], describing the experiments with arrays of Rydberg atoms [64, 97], and can be mapped onto a U(1) gauge theory [679]. In presence of disorder, $W>0$, the PXP model hosts a crossover between the ETH and MBL regimes [126]. However, the system size drifts at this crossover are significantly more pronounced than in the disordered XXZ spin chain, and both disorder strengths $W^{T}(L)$ as well as $W^{*}(L)$ shift linearly with $L$ [643].
@@
The Hilbert space dimension of the PXP model (90) scales as $\mathcal{N}\propto(\Phi_{\alpha})^{L}$, where $\Phi_{\alpha}<2$ decreases monotonously with the growth of the blockade radius $\alpha$. Hence, for larger $\alpha$ it is possible to access larger system sizes $L$ with exact diagonalization methods. Notably, at $\alpha>1$, the linear trends $W^{T}(L)\propto L$ and $W^{*}(L)\propto L$ are even more pronounced, suggesting that the PXP models (90) remain ergodic in the asymptotic limit. The nature of the kinetic constraints in $\hat{H}_{\mathrm{PXP}}$ is one factor responsible for this behavior. Indeed, the effective radius of the Fock space graph is much smaller in the PXP model than in the TFIM, enhancing the trends towards ergodicity in the former model [643]. Notably, Ref. [583] reported an opposite trend of enhancing the MBL regime due to the interplay between interaction and kinetic constraints that are of different type than those considered here.
@@
The milder growth of the Hilbert space with increase of the number of lattice sites in constrained spin models has been also used in investigations of 2D models [700, 521]. Moreover, the local constraint resemble, to some degree, the local Gauss law in lattice gauge theories. The latter have been shown to host a pronounced region in which the ergodicity is broken. The boundary of this region, i.e., the disorder strength $W^{T}(L)$, is nearly independent of system size $L$ in the 1D Schwinger model with the disordered background charges [249].